• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name reitz

    Log entry time 23:48:02 on December21,2003

    Entry number 114974

    keyword=ep target okay

    using a loose trigger, i could verify that the EP
    target is OKAY.
    needed to reboot ROC of the EP system several times, so
    i believe that the problem this morning and in the first attempt
    this evening was due to DAQ problems and NOT due to the target.
    starting another full measurements